Chapter 7 Econ Behav

19 flashcardsby shyannekendra
Study these flashcardsFree · spaced repetition · iPhone, Android, Mac & web

All cards

1
outcome space
a set of possible outcomes usually written in brackets
2
Example of outcome space
For a coin flip {H,T}
For flipping a coin twice {HH,HT,TH,TT}
3
Outcome
A subset of the outcome space
4
Example of outcome
Outcome of heads in a coin flip {H}
5
Probability Function
Is denoted as Pr().
It is a function that maps outcomes to probabilities.
6
Pr(A) = X
A is the outcome. It is the input to the probability function.
X is the probability. It is the output of the probability functno
7
Example of probability. Pr(Heads)=
0.5
8
Probabilities range from
0(will not happen with certainty) to 1(will happen with certainty)
9
The probability of rolling a seven with a six-sided die is
ZERO!
10
Everything Rule
The probability of the entire outcome set is 1
11
Example of everything rule
Rolling 1-6 with a six-sided die is 1
12
Not Rule
The probability of anything but outcome A, which we call Pr(Not A), =1-Pr(A)
13
Probability of Not Rule Example
Probability of a American roulette wheel is red.
Pr(Red) = 9/19, so the probability of any other outcome.
Pr(Not Red), is equal to 1-9/19, or 10/19
14
Equal probability
If all outcomes have equal probability, then the probability of any outcome is 1/n, where n is the number of possible outcomes
15
Independent
one occuring event DOES NOT change the probability of the other event occuring
16
Mutually Exclusive
it is impossible for both events to happen
17
Conjunction or Joint probability
Two outcomes are independent. Probability of outcome A and outcome B both happening.
Pr(A & B) is Pr(A) * Pr(B)
18
Example of Conjunction or Joint Probability
I flip a coin twice. What is the probability that the coin lands heads first and tails second (HT)?
Pr (H in first flip) = 1/2 and Pr (T in second flip) = 1/2, so Pr(HT)= 1/2 * 1/2 = 1/4
19
Disjunction
The probability of either outcome A or outcome B occuring.
Denoted Pr(A \/ B) or Pr (A or B), is Pr(A) + Pr(B) - Pr(A&B).
If the outcomes are independent, Pr (A \/ B) = Pr (A) + Pr(B) - Pr(A) * Pr (B)
If outcomes are mutually exclusive, so Pr(A&B) must equal 0, so Pr(A \/ B) = Pr(A) + Pr(B)
Study all 19 cards for free